Consider each of the following case studies. For each scenario, create a monthly budget. You will have to make reasonable assumptions for many expenses based on personal experience and internet research. Where required, include a description of any assumptions made.

Case Study 1

Part 1
Kevin has recently graduated from a college program in Computer Engineering Technology and has accepted a job with a computer networking company in Kitchener, Ontario. His monthly salary after deductions is $2235. Kevin rents a one-bedroom apartment in town. His monthly rent is $950 plus electricity, phone, internet, and cable. Kevin owns an older car which seems to need many repairs lately. Kevin is also a big movie fan. Between going to the movies and buying DVD’s, Kevin estimates that he spends $60 per week.

Part 2
Kevin borrows $12 000 to buy a new car. His car loan charges interest of 6.9% /a, compounded monthly. The loan is amortized over 3 years. Use the TVM Solver to determine his monthly payments. Create a revised budget to reflect this expense. You will likely need to reduce some of Kevin’s expenses so that he can afford his car payments.
Answer #1

Part 1:

As per internet data,

Average food expense is $550 and average electricity,phone,cable expense is $150

Average Car maintenance is $50 per month. So, below is the monthly budget for kevin:

Monthly income Rent Expense Miscellaneous Expense (Electricity, Phone, cable, internet, Food Expense) Car Maintenance Movie expense Savings (Net income-Rent-Miscellaneous Expense-Car Maintenance-Movie expense)
$2,235.00 $950.00 $700.00 $50.00 $240.00 $295.00

Part 2:

His monthly payment buying a new car would be $169.98. Below is the calculation:

Monthly interest rate=6.9%/12=0.58%

Tenure=3*12=36 month

A B C D Rate 0.58% 12000 36 PV nper (tenure) |=PMT(B1,B3,B2) PMT PMT(rate, nper, pv, [fv], [typel)

A B Rate 0.58% PV 12000 nper (tenure) PMT 36 ($369.98)

His new monthly budget is given below:

Monthly income Rent Expense Miscellaneous Expense (Electricity, Phone, cable, internet, Food Expense) Car Maintenance Movie expense Monthly Loan payment on Car Loan Savings (Net income-Rent-Miscellaneous Expense-Car Maintenance-Movie expense-CAR EMI)
$2,235.00 $950.00 $700.00 $25.00 $120.00 $369.98 $70.02

As it's a new car he has lower maintenance cost on car, I assumed this to be $25/month and reduced expense on movie s $120/month so that he can afford his car payments.